A loan of $370,000 is amortized over 30 years with payments at the end of each month and an interest rate of 8.9%, compounded monthly.



Answer the following, rounding to the nearest penny.

a) Find the amount of each payment. $
b) Find the total amount of interest paid during the first 15 payments. $

c) Find the total amount of interest paid over the life of the loan. $

d) Find the total of all payments made over 30 years. $


Suppose that payment number 8 is skipped and the interest owed for month 8 is added to the balance. Payments then resume as usual for the remainder of the 30 years.
e) Find the balance owing at the end of month 8. $
f) Find the balance remaining after the 360th payment. $
Answer #1

THE LOAN AMOUNT TAKEN : 3,70,000 USD

INTEREST RATE                   : 8.9 % COMPOUNDED MONTHLY

ANNUAL EMI PAYABLE INCLUSIVE OF INTEREST AND PRINCIPLE

MONTHLY RATE :(( 1+INTEREST RATE)/12)-1)

I.E ((1+8.9)/12)-1)= 0.713%

MONTHLY EMI = MONTHLY PRINCIPLE + INTEREST

INTEREST = PRINCIPLE * MONTHLY RATE

I.E 370000* 0.713% = 2638.21

PRINCIPLE AMOUNT = 221.56

A) EACH PAYMENT = 2859.77

B)TOTAL INTEREST PAID FOR 15 YEARS = 39401.98

C)TOTAL INTEREST PAID OVER THE LIFE OF THE LOAN = 659516.42

D)2859.77*30 = 85793.04

E)2625.25

F) 2839.52